Amy is 1/4 of the age of John. Together their total age is 40. How old is Amy and John
Answer by ewatrrr(24785)


Hi
their total age is 40
Let x and (40-x)represent John's and Amy's age respectively
Question states***
(40-x) = x/4
solving for x
4(40 - x) = x
160 = 5x
32 = x, John's age. Amy is 8. (40-32)
CHECKING our Answer***
8 = 32/4